OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)

OpenStudy (texaschic101):

we will use the slope formula slope = (y2 - y1) / (x2 - x1) and you have 2 sets of points... set 1 is (1,-5)......so x1 = 1 and y1 = -5 set 2 is (4,1)....so x2 = 4 and y2 = 1 now we sub slope = (1 - (-5) / (4 - 1) slope = (1 + 5) / 3 slope = 6/3 slope = 2